OpenStudy (anonymous):

A projectile launched upwards from ground has a range of 180m and its timeof flight is 10 s.The velocity(in m/s) after start is (angle of projection with horizontal is theta)?

OpenStudy (anonymous):

Do u know the formulae for range and time of flight?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after doing all that i am unable to get the answer......the answer should be 18

OpenStudy (anonymous):

simple \[t=\frac{2u \sin \theta}{g}=10\] range R \[R=\frac{u^2 \sin 2\theta}{g}\] use these

OpenStudy (anonymous):

@NotSObright i used this and not getting

OpenStudy (anonymous):

@NotSObright y got it.....

OpenStudy (anonymous):

Wait you are looking for both the angle and the velocity?

OpenStudy (anonymous):

no i want velocity after 5s

OpenStudy (anonymous):

also \[u \cos \theta *t\] derivable from here that \[u \cos \theta=18\]

OpenStudy (anonymous):

Well I don't think you can do it with time. http://en.wikipedia.org/wiki/Trajectory_of_a_projectile look down to where it says velocity at x not at time t find the distance traveled in 5s and plug that to the equation

OpenStudy (anonymous):

the distance after 5 seconds would be 90m or half of the total distance

OpenStudy (anonymous):

ok thanzz now i get it....

OpenStudy (anonymous):

@Brent0423 thanzz

OpenStudy (anonymous):

okay great! let me know what you come up with for the answer

OpenStudy (anonymous):

18m/s

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

If you were not provided a given starting angle then yes that answer is correct

OpenStudy (anonymous):

oh....no the angle is given as theta

OpenStudy (anonymous):

@Brent0423 is there any way

OpenStudy (anonymous):

@Brent0423 r u there

OpenStudy (anonymous):

do you know the degrees of theta? It would really help if you could draw out the problem so i can see the entire problem in order to provide you with a correct answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no angle

OpenStudy (anonymous):

can you still draw out the problem please